OpenStudy (anonymous):

The angle of elevation of a ladder leaning against a wall is 62°. If the distance from the base of the ladder to the wall is 14 feet, how long is the ladder?

OpenStudy (anonymous):

A. 14.9 feet B. 7.44 feet C. 15.87 feet D. 29.8 feet

OpenStudy (mitu12):

use cos(angle)=adjacent side/ hypotenuse

OpenStudy (aum):



OpenStudy (aum):

cos(62) = adjacent / hypotenuse = AB / AC = 14 / AC AC = 14 / cos(62) = ?

OpenStudy (anonymous):

I keep getting 6.57..

OpenStudy (aum):

set your calculator to degree mode (not radians)

OpenStudy (mitu12):

i got 29.8

OpenStudy (anonymous):

Oh okay thank you!